OpenStudy (anonymous):

how do you solve this? Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)).

OpenStudy (anonymous):

first find g(-1) then use that answer to plug it into f.

OpenStudy (anonymous):

how do i find g(-1)

OpenStudy (anonymous):

since \(\large g(x) =x^2-7x-9 \), then \(\large g(-1)=(-1)^2-7(-1)-9= \) simpify. now plug this answer into f(x) by doing the same thing with f(x)